OpenStudy (anonymous):

an some one teach me how to find the remainder of a synthetic division ___________ Heres an example 1/ 4 6 -2

OpenStudy (jdoe0001):

@Clarke11 so... what's the divisor and dividend?

OpenStudy (jdoe0001):

or is that the quotient up there? 1/ 4 6 -2

OpenStudy (anonymous):

@jdoe0001 its the division problem in synthetic form

OpenStudy (jdoe0001):

hmm? I dunno what the divisor and dividend are..... so... dunno...

OpenStudy (jdoe0001):

what are you looking for again?

OpenStudy (anonymous):

@jdoe0001 i'm looking for the remainder I think 1 is the dividend and 4 6 and -2 are the divisors but idk what synthetic form is

OpenStudy (jdoe0001):

hmmm well..... you're looking for the remainder..... in... what though? you see I dunno where we get it :).... do you have any values given to get it?

OpenStudy (anonymous):

@jdoe0001 well its multiple choice the possible answers are 9,8,6 and 7

OpenStudy (jdoe0001):

that... doesn't tell us anything though a synthetic division means there are two polynomials... and I don't see them

OpenStudy (anonymous):

@jdoe0001 i dont understand im really confused

OpenStudy (jdoe0001):

well.. how about a quick screenshot of the material?

OpenStudy (anonymous):

@jdoe0001 here ya go

OpenStudy (jdoe0001):

I see ... one sec

OpenStudy (jdoe0001):

notice that in a synthetic division you "drop down" the 1st term of the dividend, 4 in this case then multiply the divisor 1 times that, 1 * 4 and the product goes on to add "under" the next term and so on \(\large \begin{array}{lllllll} 1&|&4&6&-2\\ &|&&4&10\\ \hline\\ &&4&10&{\color{blue}{ 8}} \end{array}\qquad {\color{blue}{ remainder}}\)

OpenStudy (anonymous):

@jdoe0001 wow thats really simple, I understand how to do it but i still dont know what its used for.

OpenStudy (anonymous):

@jdoe0001 aw man theres another question the same but they're are 4 terms instead of three do i just do the same

OpenStudy (jdoe0001):

well.. the synthetic division is used to divide a polynomial by a binomial for example, your polynomials there could have been say \(\bf {\color{blue}{ 4}}x^2{\color{blue}{ +6}}x{\color{blue}{ -2}} \div x-1\qquad {\color{brown}{ x-1=0\to x={\color{blue}{ 1}}}}\)

OpenStudy (jdoe0001):

so if the binomial had been x+1 then x+1=0 => x= -1 your divisor would have been -1

OpenStudy (jdoe0001):

but that's what it's used for :)

OpenStudy (jdoe0001):

you'd do the same with whatever amount of terms, yes

OpenStudy (anonymous):

This time I have to put the quotient in polynomial form -2/2 2 -2 4 so it would be 2x^2+2x-4 right @jdoe0001

OpenStudy (jdoe0001):

well.... the quotient given, is the coefficients and the variable's degree, gets dropped by 1 for each term from left to right

OpenStudy (jdoe0001):

so if you divided by \(\bf 2x^2+2x-4\) and the quotient was say 3 4 7 that means \(\bf 3x+4\) + the remainder 7, but the 7 may need to be written in fractional form what are the actual polynomials you divided?

OpenStudy (anonymous):

@jdoe0001 Wait i have it wrong Is it 2x^2-2x+4 instead

OpenStudy (jdoe0001):

well.. is that the dividend? what about the divisor?

OpenStudy (anonymous):

@jdoe0001 here have a look

OpenStudy (jdoe0001):

\(\large \begin{array}{rrrrrrrlll} -2&|&2&2&-2&4\\ &|&&-4&4&-4 \\\hline\\ &&2&-2&2&{\color{blue}{ 0}} \end{array}\)

OpenStudy (anonymous):

@jdoe0001 I understand that, but did i put it in polynomial form correctly 2x^2+2x-4

OpenStudy (jdoe0001):

\(\large \begin{array}{rrrrrrrlll} -2&|&2&2&-2&4\\ &|&&-4&4&-4 \\\hline\\ &&2&-2&2&{\color{blue}{ 0}}\\ &&\downarrow &\downarrow &\downarrow \\ &&2x^2&-2x&+2 \end{array}\)

OpenStudy (anonymous):

@jdoe0001 I understand now thanks alot man.

OpenStudy (jdoe0001):

yw